# A Chain-of-Thought Is as Strong as Its Weakest Link: A Benchmark for Verifiers of Reasoning Chains

### v1

@@ Abstract
Prompting language models to provide step-by-step answers (e.g., “Chain-of-Thought”) is the prominent approach for complex reasoning tasks, where more accurate reasoning chains typically improve downstream task performance. Recent literature discusses automatic methods to verify reasoning steps to evaluate and improve their correctness. However, no fine-grained step-level datasets are available to enable thorough evaluation of such verification methods, hindering progress in this direction. We introduce Reveal: Reasoning Verification Evaluation, a new dataset to benchmark automatic verifiers of complex Chain-of-Thought reasoning in open-domain question answering settings. Reveal includes comprehensive labels for the relevance, attribution to evidence passages, and logical correctness of each reasoning step in a language model’s answer, across a wide variety of datasets and state-of-the-art language models.
@@ A Benchmark for Verifiers of Reasoning Chains
[…]
@@ 3.1 Task 1: Relevance, Type, and Logic
Answer: Texas population is 39.53m.${}^{(1)}$ California population is 39.24m.${}^{(2)}$ Thus, Texas has a bigger population than California.${}^{(3)}$ So the answer is Texas.${}^{(4)}$
@@ 4 Data Collection Process
We use three different LMs to generate CoT answers: Flan-PaLM-540B Chung et al. (2022), GPT-3 (text-davinci-003, Brown et al., 2020) and Flan-UL2-20B Tay et al. (2023). CoT prompt demonstrations were written according to standard practices Wei et al. (2023), with in-domain examples for each dataset taken from the dataset’s training set, and designed to be simple and informative. A subset of the questions from each dataset (one-fourth) were answered by all three models with three separate CoT answers, to serve as a base for analyses that require multiple answers per question, while the rest were answered by each model once in equal proportions (one-fourth each). Our goal in this methodology was to maximize the flexibility of the dataset and potential for analyses and evaluations, for as many LMs and verifiers as possible. Specifically, we aimed to collect a sufficient variety of answers, both correct and incorrect, from a practical and realistic distribution.
@@
Questions
704 205
CoT answers
1002 224
CoT steps
3360 847
Avg. steps per CoT
3.4 5.1
Attribution steps
1979 485
Step-evidence pairs
3502 745
Avg. evidence length (words)
103 103
Logic steps
1250 306
Fully attributable step-evidence pairs
864
Logically correct steps
1063
Fully correct CoT answers
200
@@ 5.1 Analysis of Unsupported Claims
A large fraction of the labeled steps was found unsupported by the retrieved evidence (38.6%, Figure 5). To understand this, we conduct a qualitative analysis of 20 unsupported steps (5 from each source dataset, randomly sampled). For each step, we first verify whether the step is factually correct. For correct steps, we analyzed whether the issue was due to irrelevant retrieved evidence, or due to other reasons. See Appendix A for full details.
Of the 20 unsupported steps, 8 were indeed not factual. Of the remaining 12: In 7 steps the retrieved evidence was relevant, but additional reasoning and world knowledge was required to verify them (e.g., which teams are part of the spanish “La-Liga”, or that “ACM” stands for “Association for Computing Machinery”). In 3 steps the retrieved evidence was irrelevant. In 1 step there were two claims within the given step, but only one was supported by the evidence, and in 1 step the evidence was irrelevant due to insufficient decontextualization (“Brown” ref. “Jaylen Brown”).
From this analysis, we estimate that roughly half of the unsupported cases stem from using imperfect retrieval. Additional challenges such as imperfect decontextualization and requiring to reason over multiple evidence passages can also contribute to failing to support factually correct claims.
@@ 5.2 Analysis of Reveal-Open
Complexity Category
Attribution steps (N=199)
Logic steps (N=52)
Approximating and hedging
1.51%
Averages and ranges
10.05%
Calculation or unit issue
42.31%
Specialized knowledge
13.07%
7.69%
Formatting issue
3.02%
Unclear reference
12.56%
30.77%
Inference across claims in evidence
1.51%
Invalid inference in previous step
0.50%
28.85%
Insufficient hedging
13.07%
Rating category definition/criterion
23.62%
Relevance dispute
23.08%
Temporal inconsistency
9.55%
1.92%
World knowledge
25.63%
15.38%
@@ 6.1 Step-level Verification
Given the unbalanced class distribution, we report macro-F1 performance in Table 3 (class F1 results are in Appendix B). Results for the step relevance task (classifying if a step is relevant for answering the question) are omitted, as all the models collapsed to the majority baseline, marking any step as relevant.
@@
Flan-UL2-20B
41.5 54.4
Flan-PaLM-540B
39.4 58.1
PaLM-2-L
61.9 76.4
GPT-3
35.6 71.9
@@ A.1 Source Data Collection
StrategyQA
Would a black widow have use for peaches?
A black widow is a spider. Peaches are a fruit. Thus, a black widow would have use for peaches. So the answer is yes.
Logic (step 3)
StrategyQA
Will more people go in and out of Taco Bell than a Roy Rogers each year?
Taco Bell is a fast food chain with thousands of locations. Roy Rogers is a fast food chain with fewer than 500 locations. Thus, more people will go in and out of Taco Bell than Roy Rogers each year. So the answer is yes.
Correct
MuSiQue
Who is the other child of Chian Hsiao-chang’s father?
Chian Hsiao-chang’s father is Chian Ching-kuo. Chian Hsiao-chang’s other child is Chian Ching-wen. So the final answer is: Chian Ching-wen.
Attribution (step 2), logic (step 3)
MuSiQue
Which New York City team does not host NBA games at the site where Jimmy Carter was nominated for president?
Jimmy Carter was nominated for president at Madison Square Garden in New York City. Madison Square Garden is home to the New York Knicks, but not the Brooklyn Nets. So the final answer is: Brooklyn Nets.
Correct
Sports
Is the following sentence plausible? “Jamison Crowder changed direction in the backfield.”
Jamison Crowder is a baseball player. Changing direction in the backfield is part of American football. So the answer is no.
Attribution (step 1)
Sports
Is the following sentence plausible? “Yoan Moncada beat the buzzer.”
Yoan Moncada is a baseball player. Beating out the buzzer is part of basketball. So the answer is no.
Correct
Fermi
What is the volume, cc, of blood in a healthy new-born baby?
The mass of blood in a new born baby is 25 grams. The density of blood is 0.994 g*cm**-3 or grams per cc. Thus, the cc volume of blood in a new-born baby is 25g / 0.994 g*cm**-3 = 25.15 cm**3. So the final answer is: 25.15
Attribution (step 1)
Fermi
What is the total US salary?
The per capita income in the US is 63051 USD. There are around 123e+6 full time employees in the US. Thus, the total US salary is \$ 63051 * 123e+6 = 7755273000000 USD. So the final answer is: 7755273000000
Correct
@@ B.3 Prompt Templates
Attribution (N=199)
Logic (N=52)
World knowledge or general inference
25.63%
15.38%
Some annotators did not apply world knowledge or general implicit inferences that other annotators take for granted (e.g., that a civil war takes place within a single country, or that an animal described as having prey can be called a “predator”).
Acceptable nuance/hedging
1.51%
Disagreement between annotators on strictness of hedged approximations in the step. E.g., “There are around 7.5e+9 people in the world” where the evidence mentions “7.55 billion”.
Insufficient nuance/hedging
13.07%
A step mentions an approximation of some quantity in the evidence but without appropriate hedging.
Averages and ranges
10.05%
A step picks a specific point or subrange from a more general range or set of ranges given in the evidence, and annotators disagree on whether the exemplification is representative of the evidence.
Calculation or unit issue
42.31%
The step consists of a calculation or unit conversion which is difficult to verify.
Specialized knowledge
13.07%
7.69%
The step requires specialized knowledge, e.g., rules of a sports game or scientific notation, which some annotations consider as world knowledge.
Formatting issue
3.02%
A problem in the annotation interface (e.g., truncation) made the claims impossible to verify.
Inconsistent/unclear reference or standard
12.56%
30.77%
A step makes an ambiguous reference. E.g., “Westworld was directed by Jonathan Nolan” is paired with evidence on “Westworld TV series …with Nolan directing the pilot episode”—Westworld could refer to the film, TV series, etc.
Inference across claims in evidence
1.51%
Verifying the step requires drawing an implicit conclusion, which only some annotators did—e.g., the step “Birds and mammals are different classes of animals” is paired with evidence that describes characteristics of mammals, and mentions that they “distinguish them from reptiles and birds”, which implies that they are in different classes.
Invalid inference in previous step
0.50%
28.85%
Some annotators may have found it difficult to disentangle a current step’s inference from any previous inferences that are invalid.
Rating category definition/criterion
23.62%
Disagreements stemming from the distinction between “partially attributable” and “not attributable” (despite guidance during the annotation instructions—more discussion below).
Relevance dispute
23.08%
Borderline-relevant steps or difficult to follow steps can lead to disagreement on the relevance label.
Temporal inconsistency
9.55%
1.92%
The step makes a claim that relies on a different time frame from that of the evidence or other steps in the answer (e.g., present-tense in “Maradona is a soccer player” where the evidence mentions “a former professional footballer”). Annotators’ acceptance of temporal inconsistencies may vary (e.g., the convention of using the present tense to refer to a retired person’s most salient profession).
@@
Fermi
There are 1000 * 100 = 10000 pennies in 1 US dollar.
No
Wrong math
Fermi
Art Beat is a free, annual event in South Bend, Indiana.
Yes
@@
The info could be found outside of Wikipedia, the retrieved evidence was irrelevant.
Fermi
The latent heat of vaporization of steam is 540 cal/g.
Yes
@@
The info could be found outside of wikipedia, the retrieved evidence was irrelevant.
Fermi
The total volume of the oceans is 1.3e+21 liters.
No
@@
The total volume of the oceans is much more than 1.3e+12 liters. The retrieved evidence was relevant but did not mention the answer explicitly (metric conversion was required).
Fermi
The average American yard is 8000 square feet.
No
@@
Could not find evidence that mentions this number, and the retrieved evidence was irrelevant.
Musique
Jayantha Ketagoda is married to his wife, Jaya.
No
@@
Could not find evidence that mentions this claim, and the retrieved evidence was relevant to the entity but did not mention the claim.
Musique
@@
Vibullia Alcia Agrippina’s child was a proponent for the movement called the ”Feminist Movement”.
No
Could not find evidence that mentions this claim.
Musique
The Almost are signed to the record label called Almost Records.
No
@@
The claim is wrong, and the retrieved evidence was not relevant. Could find evidence in wikipedia that refutes the claim.
Musique
The most Champions League wins between 1992 and 2013 were by the Spanish La Liga.
Yes
@@
The claim is correct, but requires complex reasoning (understanding which teams are from the Spanish La Liga). The retrieved evidence was related to La Liga but not sufficient to support the claim.
Musique
Deborah Estrin is a member of the Association for Computing Machinery.
Yes
@@
The claim is correct. There was evidence that supported the claim, but this required domain knowledge (that “ACM”=“Association for Computing Machinery”) so the raters marked it as unsupported.
Sports
Doing a maradona on the defender is part of basketball.
No
@@
The claim is wrong, “a maradona” is a soccer move, not a basketball move.
Sports
Hitting the back of the rim is a common occurrence in basketball.
Yes
@@
The claim is correct. None of the evidence supported it directly, as it is more “common knowledge”.
Sports
Windmill dunk is part of basketball, not hockey.
Yes
@@
The claim is correct, however it is a composite claim - 1) Windmill dunk is part of basketball 2) Windmill dunk is not part of hockey. The evidence supported the first claim, not the second.
Sports
Brown is a basketball player.
Yes
@@
The claim is correct, but requires further decontextualization (from “Brown” to “Jaylen Brown”) as it is vague in its current form, which resulted in irrelevant evidence.
Sports
Being out at second is part of baseball.
Yes
@@
The claim is correct, but requires common knowledge - that “out at second” means “second base”. The retrieved evidence did not support this claim explicitly.
StrategyQA
A king size bed is 76 inches wide and 80 inches long.
Yes
@@
The step is factually correct and a google search was able to retrieve relevant evidence. The retrieved evidence was irrelevant.
StrategyQA
Pacifists do not support violence, including stoning.
Yes
@@
The step is factually correct, and can be verified with common sense. The retrieved evidence did not support this claim.
StrategyQA
An existential crisis is a mental health issue.
Yes
@@
The step is factually correct, and can be verified with common sense. The retrieved evidence did not support this claim explicitly.
StrategyQA
Lil Jon’s top ranked Billboard song was "Get Low".
No
@@
The claim is wrong since the song is not the only top-ranked song on Billboard from the artist. The retrieved evidence was relevant and partially supported the claim (saying that the song was indeed top-ranked in the chart).
StrategyQA
Cricketers do not kick field goals.
Yes
@@
The claim is correct since a field goal is a term from other sports and kicking is generally prohibited in Cricket. The retrieved evidence was irrelevant since this requires more common-knowledge and inference.

### v4

@@ Abstract
Prompting language models to provide step-by-step answers (e.g., “Chain-of-Thought”) is the prominent approach for complex reasoning tasks, where more accurate reasoning chains typically improve downstream task performance. Recent literature discusses automatic methods to verify reasoning steps to evaluate and improve their correctness. However, no fine-grained step-level datasets are available to enable thorough evaluation of such verification methods, hindering progress in this direction. We introduce Reveal: Reasoning Verification Evaluation, a new dataset to benchmark automatic verifiers of complex Chain-of-Thought reasoning in open-domain question answering settings. Reveal includes comprehensive labels for the relevance, attribution to evidence passages, and logical correctness of each reasoning step in a language model’s answer, across a wide variety of datasets and state-of-the-art language models. Available at reveal-dataset.github.io.
@@ A Benchmark for Verifiers of Reasoning Chains
Alon Jacovi1,2  Yonatan Bitton1  Bernd Bohnet3  Jonathan Herzig1 Or Honovich1,4  Michael Tseng3  Michael Collins3  Roee Aharoni1  Mor Geva1,4 1Google Research  2Bar Ilan University  3Google DeepMind  4Tel Aviv University alonjacovi@google.com
@@ 3.1 Task 1: Relevance, Type, and Logic
Answer: Texas population is 39.53m.(1) California population is 39.24m.(2) Thus, Texas has a bigger population than California.(3) So the answer is Texas.(4)
@@ 4 Data Collection Process
We use three different LMs to generate CoT answers: Flan-PaLM-540B Chung et al. (2022), GPT-3 (text-davinci-003, Brown et al., 2020) and Flan-UL2-20B Tay et al. (2023). We prioritized a variety in the models, with two large-size high-performing models with different pretraining data, and one smaller-size, weaker model, in order to gather a variety of CoT answers. CoT prompt demonstrations were written according to standard practices Wei et al. (2023), with in-domain examples for each dataset taken from the dataset’s training set, and designed to be simple and informative. A subset of the questions from each dataset (one-fourth) were answered by all three models with three separate CoT answers, to serve as a base for analyses that require multiple answers per question, while the rest were answered by each model once in equal proportions (one-fourth each). Our goal in this methodology was to maximize the flexibility of the dataset and potential for analyses and evaluations, for as many LMs and verifiers as possible. Specifically, we aimed to collect a sufficient variety of answers, both correct and incorrect, from a practical and realistic distribution.
@@
Questions 704 205
CoT answers 1002 224
CoT steps 3360 847
Avg. steps per CoT 3.4 5.1
Attribution steps 1979 485
Step-evidence pairs 3502 745
Avg. evidence length (words) 103 103
Logic steps 1250 306
Fully attributable step-evidence pairs 864
Logically correct steps 1063
Fully correct CoT answers 200
@@ 5.1 Analysis of Unsupported Claims
A large fraction of the labeled steps was found unsupported by the retrieved evidence (38.6%, Figure 5). To understand this, we conduct a qualitative analysis of 40 unsupported steps (10 from each source dataset, randomly sampled). For each step, we first verify whether the step is factually correct. For correct steps, we analyzed whether the issue was due to irrelevant retrieved evidence, or due to other reasons. See Appendix A for full details.
Of the 40 unsupported steps, 19 were indeed not factual. Of the remaining 21: In 13 steps the retrieved evidence was relevant, but additional reasoning and world knowledge was required to verify them (e.g., which teams are part of the spanish “La-Liga”, or that “ACM” stands for “Association for Computing Machinery”). In 6 steps the retrieved evidence was irrelevant. In 1 step there were two claims within the given step, but only one was supported by the evidence, and in 1 step the evidence was irrelevant due to insufficient decontextualization (“Brown” ref. “Jaylen Brown”).
From this analysis, we estimate that roughly half of the unsupported cases stem from using imperfect retrieval. Additional challenges such as imperfect decontextualization and requiring to reason over multiple evidence passages can also slightly contribute to failing to support factually correct claims.
@@ 5.2 Analysis of Reveal-Open
Complexity Category Attribution steps (N=199) Logic steps (N=52)
Approximating and hedging 1.51%
Averages and ranges 10.05%
Calculation or unit issue 42.31%
Specialized knowledge 13.07% 7.69%
Formatting issue 3.02%
Unclear reference 12.56% 30.77%
Inference across claims in evidence 1.51%
Invalid inference in previous step 0.50% 28.85%
Insufficient hedging 13.07%
Rating category definition/criterion 23.62%
Relevance dispute 23.08%
Temporal inconsistency 9.55% 1.92%
World knowledge 25.63% 15.38%
@@ 6.1 Step-level Verification
Given the unbalanced class distribution, we report macro-F1 performance in Table 3 (class F1 results are in Appendix B). Results for the step relevance task (classifying if a step is relevant for answering the question) are omitted, as all the models collapsed to the majority baseline, leading to an F1 of near-0.
@@
Flan-UL2-20B 41.5 54.4
Flan-PaLM-540B 39.4 58.1
PaLM-2-L 61.9 76.4
GPT-3 35.6 71.9
@@ A.1 Source Data Collection
StrategyQA Would a black widow have use for peaches? A black widow is a spider. Peaches are a fruit. Thus, a black widow would have use for peaches. So the answer is yes. Logic (step 3)
StrategyQA Will more people go in and out of Taco Bell than a Roy Rogers each year? Taco Bell is a fast food chain with thousands of locations. Roy Rogers is a fast food chain with fewer than 500 locations. Thus, more people will go in and out of Taco Bell than Roy Rogers each year. So the answer is yes. Correct
MuSiQue Who is the other child of Chian Hsiao-chang’s father? Chian Hsiao-chang’s father is Chian Ching-kuo. Chian Hsiao-chang’s other child is Chian Ching-wen. So the final answer is: Chian Ching-wen. Attribution (step 2), logic (step 3)
MuSiQue Which New York City team does not host NBA games at the site where Jimmy Carter was nominated for president? Jimmy Carter was nominated for president at Madison Square Garden in New York City. Madison Square Garden is home to the New York Knicks, but not the Brooklyn Nets. So the final answer is: Brooklyn Nets. Correct
Sports Is the following sentence plausible? “Jamison Crowder changed direction in the backfield.” Jamison Crowder is a baseball player. Changing direction in the backfield is part of American football. So the answer is no. Attribution (step 1)
Sports Is the following sentence plausible? “Yoan Moncada beat the buzzer.” Yoan Moncada is a baseball player. Beating out the buzzer is part of basketball. So the answer is no. Correct
Fermi What is the volume, cc, of blood in a healthy new-born baby? The mass of blood in a new born baby is 25 grams. The density of blood is 0.994 g*cm**-3 or grams per cc. Thus, the cc volume of blood in a new-born baby is 25g / 0.994 g*cm**-3 = 25.15 cm**3. So the final answer is: 25.15 Attribution (step 1)
Fermi What is the total US salary? The per capita income in the US is 63051 USD. There are around 123e+6 full time employees in the US. Thus, the total US salary is \$ 63051 * 123e+6 = 7755273000000 USD. So the final answer is: 7755273000000 Correct
@@ A.3 Additional Statistics
Each step’s annotation took between 300 to 600 seconds, with some answers having 5 or more steps, and with 5 annotations per step—leading to an expensive annotation process.
@@ B.3 Prompt Templates
Attribution (N=199) Logic (N=52)
World knowledge or general inference 25.63% 15.38% Some annotators did not apply world knowledge or general implicit inferences that other annotators take for granted (e.g., that a civil war takes place within a single country, or that an animal described as having prey can be called a “predator”).
Acceptable nuance/hedging 1.51% Disagreement between annotators on strictness of hedged approximations in the step. E.g., “There are around 7.5e+9 people in the world” where the evidence mentions “7.55 billion”.
Insufficient nuance/hedging 13.07% A step mentions an approximation of some quantity in the evidence but without appropriate hedging.
Averages and ranges 10.05% A step picks a specific point or subrange from a more general range or set of ranges given in the evidence, and annotators disagree on whether the exemplification is representative of the evidence.
Calculation or unit issue 42.31% The step consists of a calculation or unit conversion which is difficult to verify.
Specialized knowledge 13.07% 7.69% The step requires specialized knowledge, e.g., rules of a sports game or scientific notation, which some annotations consider as world knowledge.
Formatting issue 3.02% A problem in the annotation interface (e.g., truncation) made the claims impossible to verify.
Inconsistent/unclear reference or standard 12.56% 30.77% A step makes an ambiguous reference. E.g., “Westworld was directed by Jonathan Nolan” is paired with evidence on “Westworld TV series …with Nolan directing the pilot episode”—Westworld could refer to the film, TV series, etc.
Inference across claims in evidence 1.51% Verifying the step requires drawing an implicit conclusion, which only some annotators did—e.g., the step “Birds and mammals are different classes of animals” is paired with evidence that describes characteristics of mammals, and mentions that they “distinguish them from reptiles and birds”, which implies that they are in different classes.
Invalid inference in previous step 0.50% 28.85% Some annotators may have found it difficult to disentangle a current step’s inference from any previous inferences that are invalid.
Rating category definition/criterion 23.62% Disagreements stemming from the distinction between “partially attributable” and “not attributable” (despite guidance during the annotation instructions—more discussion below).
Relevance dispute 23.08% Borderline-relevant steps or difficult to follow steps can lead to disagreement on the relevance label.
Temporal inconsistency 9.55% 1.92% The step makes a claim that relies on a different time frame from that of the evidence or other steps in the answer (e.g., present-tense in “Maradona is a soccer player” where the evidence mentions “a former professional footballer”). Annotators’ acceptance of temporal inconsistencies may vary (e.g., the convention of using the present tense to refer to a retired person’s most salient profession).
@@
Fermi There are 1000 * 100 = 10000 pennies in 1 US dollar. No Wrong math
@@
Fermi Art Beat is a free, annual event in South Bend, Indiana. Yes The info could be found outside of Wikipedia, the retrieved evidence was irrelevant.
@@
Fermi The latent heat of vaporization of steam is 540 cal/g. Yes The info could be found outside of wikipedia, the retrieved evidence was irrelevant.
@@
Fermi The total volume of the oceans is 1.3e+21 liters. No The total volume of the oceans is much more than 1.3e+12 liters. The retrieved evidence was relevant but did not mention the answer explicitly (metric conversion was required).
@@
Fermi The average American yard is 8000 square feet. No Could not find evidence that mentions this number, and the retrieved evidence was irrelevant.
@@
Musique Jayantha Ketagoda is married to his wife, Jaya. No Could not find evidence that mentions this claim, and the retrieved evidence was relevant to the entity but did not mention the claim.
@@
Musique Vibullia Alcia Agrippina’s child was a proponent for the movement called the ”Feminist Movement”. No Could not find evidence that mentions this claim.
@@
Musique The Almost are signed to the record label called Almost Records. No The claim is wrong, and the retrieved evidence was not relevant. Could find evidence in wikipedia that refutes the claim.
@@
Musique The most Champions League wins between 1992 and 2013 were by the Spanish La Liga. Yes The claim is correct, but requires complex reasoning (understanding which teams are from the Spanish La Liga). The retrieved evidence was related to La Liga but not sufficient to support the claim.
@@
Musique Deborah Estrin is a member of the Association for Computing Machinery. Yes The claim is correct. There was evidence that supported the claim, but this required domain knowledge (that “ACM”=“Association for Computing Machinery”) so the raters marked it as unsupported.
@@
Sports Doing a maradona on the defender is part of basketball. No The claim is wrong, “a maradona” is a soccer move, not a basketball move.
@@
Sports Hitting the back of the rim is a common occurrence in basketball. Yes The claim is correct. None of the evidence supported it directly, as it is more “common knowledge”.
@@
Sports Windmill dunk is part of basketball, not hockey. Yes The claim is correct, however it is a composite claim - 1) Windmill dunk is part of basketball 2) Windmill dunk is not part of hockey. The evidence supported the first claim, not the second.
@@
Sports Brown is a basketball player. Yes The claim is correct, but requires further decontextualization (from “Brown” to “Jaylen Brown”) as it is vague in its current form, which resulted in irrelevant evidence.
@@
Sports Being out at second is part of baseball. Yes The claim is correct, but requires common knowledge - that “out at second” means “second base”. The retrieved evidence did not support this claim explicitly.
@@
StrategyQA A king size bed is 76 inches wide and 80 inches long. Yes The step is factually correct and a google search was able to retrieve relevant evidence. The retrieved evidence was irrelevant.
@@
StrategyQA Pacifists do not support violence, including stoning. Yes The step is factually correct, and can be verified with common sense. The retrieved evidence did not support this claim.
@@
StrategyQA An existential crisis is a mental health issue. Yes The step is factually correct, and can be verified with common sense. The retrieved evidence did not support this claim explicitly.
@@
StrategyQA Lil Jon’s top ranked Billboard song was "Get Low". No The claim is wrong since the song is not the only top-ranked song on Billboard from the artist. The retrieved evidence was relevant and partially supported the claim (saying that the song was indeed top-ranked in the chart).
@@
StrategyQA Cricketers do not kick field goals. Yes The claim is correct since a field goal is a term from other sports and kicking is generally prohibited in Cricket. The retrieved evidence was irrelevant since this requires more common-knowledge and inference.